OpenStudy (anonymous):

Help me solve for x in: 14-(6x-5)+5(2x-1)-4x step by step so I understand, thank you!

OpenStudy (anonymous):

did you get my message?

OpenStudy (anonymous):

14-(6x-5)=5(2x-1)-4x First, distribute to get rid of parenthesis 14 - 6x + 5 = 10x -5 - 4x Next, combine like terms on each side and last, get the x all by itself. so you combine terms and tell me what you got

OpenStudy (anonymous):

at the end, yes you will put x on one side and numbers on the other. But first, combine the like terms on each side before you start moving things around...

OpenStudy (anonymous):

I did, and thanks

OpenStudy (anonymous):

I did this: 14-6x+5=10x-5-4x then I added six on both sides 14-5=16x-5-4x then I added 5 on both sides 14=16x-10-4x then I made the 16x- -4x and got -20x so i did 14=-20x - 10 then made it =10 and put it on the other side 14+10=24 24=-20x I don't know how I can divide that.